OpenStudy (anonymous):

Suppose you have 30 CDs. You know that you have 11 more CDs than your friend. Write and solve an equation to find the number of CDs your friend has.

Miracrown (miracrown):

Let's identify some variables. Let's let y = number of CDs you have more than your friend. And let's let x be the number of CD's that your friend has.

Miracrown (miracrown):

So, we know that you have 30 CDs, right? And your friend has x CDs And you want to know the difference between them so that you can calculate y which is how many more CDs you have than your friend. What then would be the formula for this problem?

OpenStudy (anonymous):

30 = 11 + n Subtract 11 from both sides of the equation: 30-11 = 11 - 11 + n and you get: n = 19. ?

Miracrown (miracrown):

Nope, that is incorrect ^

Miracrown (miracrown):

You put the numbers in the wrong spot. The equation is y = 30 - x

Miracrown (miracrown):

Now, 11 is the y because 11 is how many CDs you have more than your friend and y is defined as how many CDs you have more than your friend. And x is what we are solving for.

Miracrown (miracrown):

That is the what we have to solve. After plugging in y = 11 we get. 11 = 30 - x Subtracting 30 on both sides gets us -19 = -x Dividing both sides by -1 gets us: 19 = x And therefore we arrive at our answer of x = 19 which means that our friend has 19 CDs.

OpenStudy (anonymous):

ohhhhh ok awesome I finally got it! lol thank you!!